# Set Features for Anomaly Detection

### v3

@@ 8 Conclusion
9 Acknowledgement
Niv Cohen was funded by Israeli Science Foundation and the Hebrew University Data Science grants (CIDR). We thank Paul Bergmann for kindly sharing numerical results for many of the methods compared on the MVTec-LOCO dataset.
Supplementary Material
10 Full Results Tables
The full table image logical anomaly detection experiments can be found in Tab.6. The full table for the time series anomaly detection experiments can be found in Tab.7.
@@ 10 Full Results Tables
Logical Anomalies Breakfast box 70.3 58.0 47.3 69.4 59.9
@@
Logical Anomalies Breakfast box 68.9 81.8 77.7 87.0 97.7 $\pm$ 0.2
@@
[…]
OCSVM IF LOF RNN ED
EPSY 61.1 67.7 56.1 80.4 82.6
NAT 86.0 85.4 89.2 89.5 91.5
SAD 95.3 88.2 98.3 81.5 93.1
CT 97.4 94.3 97.8 96.3 79.0
RS 70.0 69.3 57.4 84.7 65.4
Avg. 82.0 81.0 79.8 86.5 82.3
DSVDD DAGMM GOAD DROCC NeuTraL Ours
EPSY 57.6 72.2 76.7 85.8 92.6 98.1
NAT 88.6 78.9 87.1 87.2 94.5 96.1
SAD 86.0 80.9 94.7 85.8 98.9 97.8
CT 95.7 89.8 97.7 95.3 99.3 99.7
RS 77.4 51.0 79.9 80.0 86.5 92.3
Avg. 81.1 74.6 87.2 86.8 94.4 96.8
11 UEA Results with Standard Errors
We present an extended version of the UEA results including error bounds for our method and baselines that reported them. The difference between the methods is significantly larger than the standard error.
@@ 11 UEA Results with Standard Errors
12 Set descriptor comparison
Clustering-based set descriptors. We compare our histogram-based approach to the VLAD and Bag-of-Features approaches. It can be seen that while effective, they still underperform our method. We do not report the results on Fisher-Vectors as the underlying GMM model (unlike K-means) requires unfeasible computational resources with our set dimensions. Taken together, it seems that the underlying clustering assumption does not fit the sets we wish to describe as well our set descriptors (we report in Tab.9 the results for $C=100$ cluster, but this result persists when we varied the number of clusters).
$k$NN versus distance to the mean. We found that using the Gaussian model only to whiten the data and taking the distance to the $1$ nearest neighbors worked better for the MVTec-LOCO dataset (see Tab.9). The nearest neighbors density estimation algorithm better models the density distribution when the Gaussian assumption is not an accurate description of the data.
@@ 12 Set descriptor comparison
13 Implementation Details
Histograms. In practice, we use the cumulative histograms as our set features for both data modalities (of Sec.3.3).
13.1 Image anomaly detection
ResNet levels. We use the representations from the 3rd and 4th blocks of a WideResNet50$\times$2 (resulting in sets size $7\times 7$ and $14\times 14$ elements, respectively). We also use all the raw pixels in the image as an additional set (resized to $224\times 224$ elements). The total anomaly score is the average of the anomaly scores obtained for the set of 3rd ResNet block features, the set of 4th ResNet block features, and the set of raw pixels. The average anomaly score is weighted by the following factors $(1,1,0.1)$ respectively (see App.14 for our robustness to the choice of weighting factor).
Parameters. For the image experiments, we use histograms of $K=5$ bins and $r=1000$ projections. For the raw-pixels layer, we used a projection dimension of $r=10$ and no whitening due to the low number of channels. To avoid high variance between runs, we did $32$ different repetitions for the raw-pixel scoring and used the median. We use $k=1$ for the $k$NN density estimation.
@@ 13.1 Image anomaly detection
Software. For the whitening of image features we use the ShrunkCovariance function from the scikit-learn library [39] with its default parameters. For $k$NN density estimation we use the faiss library [31].
@@
13.2 Time Series anomaly detection
Padding. Prior to window extraction, the series $x$ is first right and left zero-padded by $\frac{\tau}{2}$ to form padded series $x^{\prime}$. The first window $w_{1}$ is defined as the first $\tau$ observations in padded series $S^{\prime}$, i.e. $w_{1}=x^{\prime}_{1},x^{\prime}_{2}..x^{\prime}_{\tau}$. We further define windows at higher scales $W^{s}$, which include observations sampled with stride $c$. At scale $c$, the original series $x$ is right and left zero-padded by $\frac{c\cdot\tau}{2}$ to form padded series $S^{\prime c}$.
UEA Experiments. We used each time series as an individual training sample. We chose a kernel size of $9$, $100$ projection, $20$ quantiles, and a maximal number of levels of $10$. The results varied only slightly within a reasonable range of the hyperparameters e.g. using $5$, $10$, $15$ levels yielded an average ROCAUC of $97$, $96.8$, $96.8$ across the five UEA datasets.
Spoken Arabic Digits processing We follow the processing of the dataset as done by Qiu et al. [42]. In private communications the authors explained that only sequences of lengths between 20 and 50 time steps were selected. The other time series were dropped.
Computational resources. The experiments were run on a modest number of CPUs on a computing cluster. The baseline methods were run on a single RTX2080-GT GPU
13.3 License:
The package faiss [31] used for $k$NN ”MIT License”.
14 Logical Anomaly Detection Robustness
We check the robustness of our results for the parameter $\lambda$ the weighting between the raw-pixels level anomaly score to the anomaly score derived from the ResNet features (Sec.5.3). As can be seen in Tab.10, our results are robust to the choice of $\lambda$.
@@ 14 Logical Anomaly Detection Robustness
15 Further Image Anomaly Detection Ablation
We include here the ablation tables for the number of random projections and number of bins, for logical anomaly detection (Tab.11,12).
@@ 15 Further Image Anomaly Detection Ablation
16 Time Series Anomaly Detection Ablations
@@
Number of bins. We compute the accuracy of our method as a function of the number of bins per projection. Our results ( Fig. 4) show that beyond a very small number of bins - larger numbers are not critical. We found $20$ bins to be sufficient in all our experiments.
Figure 4: Ablation of accuracy vs. the number of projections (left) and the number of bins (right).
@@
No whitening 62.1 70.9 93.6 98.5 78.8
Whitening 98.1 92.3 96.1 99.7 97.8
Effect of Gaussian density estimation. Standard projection methods such as HBOS [23] and LODA [41] do not use a multivariate density estimator but instead estimate the density of each dimension independently. We compare using a full and per-variable density estimation in Tab. 13. We can see that our approach achieves far better results, attesting to the importance of modeling the correlation between projections.
@@
Effect of number of pyramid levels and window size. We ablate the two hyperparameters of the time-series feature extraction: the number of pyramid windows used $L$, and the number of samples per window $\tau$ (see Sec.4.2). We find that in both cases the results are not sensitive to the chosen parameters (Tab.15,16).
@@
17 Using the Central Limit Theorem for Set Anomaly Detection
We model the features of each window $f$ as a normal set as IID observations coming from a probability distribution function $p(f)$. The distribution function is not assumed to be Gaussian. Using a Gaussian density estimator trained on the features of elements observed in training is unlikely to be effective for element-level anomaly detection (due to the non-Gaussian $p(f)$).
An alternative formulation to the one presented in Sec. 3, is that each feature $f$ is multiplied by projection matrix $P$, and then each dimension is discretized and mapped to a one-hot vector. This formulation therefore maps the representation of each element to a sparse binary vector. The mean of the representations of elements in the set recovers the normalized histogram descriptor precisely (therefore this formulation is equivalent to the one in Sec. 3). As the histogram is a mean of the set of elements, it has superior statistical properties. In particular, the Central Limit Theorem states that under some conditions the sample mean follows the Gaussian distribution regardless of the distribution of windows $p(f)$. While typically in anomaly detection only a single sample is presented at a time, the situation is different when treating samples as sets. Although the windows are often not IID, given a multitude of elements, an IID approximation may be approximately correct. This explains the high effectiveness of Gaussian density estimation in our formulation.

### v4

@@ 8 Conclusion
9 Acknowledgment
This research was partially supported by the Israeli Science Foundation, the Israeli Council for Higher Education (VATAT), the Hebrew University Data Science grants (CIDR), the Malvina and Solomon Pollack scholarship, and a grant from KLA. Niv Cohen was partially supported by the Israeli data science scholarship for outstanding postdoctoral fellows (VATAT). We thank Paul Bergmann for kindly sharing numerical results for many of the methods compared on the MVTec-LOCO dataset.
Appendix A Full Results Tables
The full table for the image logical anomaly detection experiment can be found in Tab.6. The full table for the time series anomaly detection experiments including error bounds for our method and baselines that reported them can be found in Tab.7. The difference between the methods is significantly larger than the standard error.
@@ Appendix A Full Results Tables
Logical Anom. Breakfast box 70.3 58.0 47.3 69.4 59.9
@@
Logical Anom. Breakfast box 68.9 81.8 77.7 87.0 97.7 $\pm$ 0.2
@@
Appendix B Set descriptor comparison
Clustering-based set descriptors. We compare our histogram-based approach to the VLAD and Bag-of-Features approaches. We report our results in Tab.8. We use the number of means $K=100$ cluster, but this result persists when we varied the number of clusters. We do not report the results on Fisher-Vectors as the underlying GMM model (unlike K-means) requires unfeasible computational resources with our set dimensions. Taken together, it seems that the underlying clustering assumption does not fit the sets we wish to describe as well our set descriptors.
$k$NN versus distance to the mean. We found that using the Gaussian model only to whiten the data and taking the distance to the $1$ nearest neighbors (and not to the center) worked better for the MVTec-LOCO dataset (see Tab.8). The nearest neighbors density estimation algorithm better models the density distribution when the Gaussian assumption is not an accurate description of the data.
@@ Appendix B Set descriptor comparison
Appendix C Implementation Details
Histograms. We use the cumulative histograms as our set features for both data modalities (of Sec.3.3).
C.1 Image Anomaly Detection
Parameters. For the image experiments, we use histograms of $K=5$ bins and $r=1000$ projections. For the raw-pixels layer, we used a projection dimension of $r=10$ and no whitening due to the low number of channels. To avoid high variance between runs, we did $32$ different repetitions for the raw-pixel scoring and used the median score. We use $k=1$ for the $k$NN density estimation.
@@ C.1 Image Anomaly Detection
Software. For the whitening of image features we use the ShrunkCovariance function from the scikit-learn library Pedregosa et al. (2011) with its default parameters. For $k$NN density estimation we use the faiss library Johnson et al. (2019).
@@
C.2 Time Series Anomaly Detection
UEA Experiments. We used each time series as an individual training sample. We chose a kernel size of $\tau=9$, $r=100$ projection, $K=20$ quantiles, and a maximal number of pyramid levels of $L=10$, each using consecutive strides. The results varied only slightly within a reasonable range of the hyperparameters. E.g. using $5$, $10$, $15$ levels yielded an average ROCAUC of $97$, $96.8$, $96.8$ across the five UEA datasets.
Padding. Prior to window extraction, the series $x$ is first right and left zero-padded by $\frac{\tau}{2}$ to form a padded series $x^{\prime}$. The first window $w_{1}$ is defined as the first $\tau$ observations in padded series $S^{\prime}$, i.e. $w_{1}=x^{\prime}_{1},x^{\prime}_{2}..x^{\prime}_{\tau}$. We further define windows at higher scales $W^{s}$, which include observations sampled with stride $c$. At scale $c$, the original series $x$ is right and left zero-padded by $\frac{c\cdot\tau}{2}$ to form padded series $S^{\prime c}$.
Spoken Arabic Digits processing We follow the processing of the dataset as done by Qiu et al. Qiu et al. (2021). In private communications the authors explained that only use sequences of lengths between 20 and 50 time steps were selected. The other time series were dropped.
Computational resources. The experiments were run on a modest number of CPUs on a computing cluster. The baseline methods were run on a single RTX2080-GT GPU.
Appendix D Logical Anomaly Detection Robustness
We check the robustness of our results to the parameter $\lambda$ the weighting between the raw-pixels level anomaly score to the anomaly score derived from the ResNet features (Sec.5.3). As can be seen in Tab.9, our results are robust to the choice of $\lambda$.
@@ Appendix D Logical Anomaly Detection Robustness
Appendix E Further Image Anomaly Detection Ablation
We include here the ablation tables for the number of random projections and number of bins, for logical anomaly detection (Tab.10,11).
@@ Appendix E Further Image Anomaly Detection Ablation
Appendix F Time Series Anomaly Detection Ablations
@@
Number of bins. We compute the accuracy of our method as a function of the number of bins per projection. Our results (Fig. 4) show that beyond a very small number of bins, a larger number of bins does not help. We found $20$ bins to be sufficient in all our experiments.
Figure 4: Ablation of accuracy vs. the number of projections (left) and the number of bins (right) for different time series datasets.
@@
Quantized SWD (No whitening) 62.1 70.9 93.6 98.5 78.8
SWD 90.7 84.8 91.9 99.3 88.4
With Whitening 98.1 92.3 96.1 99.7 97.8
Effect of Gaussian density estimation. We ablate the use of Gaussian modeling in Tab. 12 (Quantized SWD). We can see that our approach achieves far better results, attesting to the importance of modeling the correlation between projections.
Comparison to the Sliced Wasserstein distance. In Sec.3.5 we highlight the connection between our approach and the Sliced Wasserstein distance. An empirical comparison between the approaches can be found in Tab. 12. Our results show that computing the SWD without histogram binning can be much more accurate than with binning (Quantized SWD). However, the binning is necessary for our whitening technique (Sec.3.4), which significantly outperforms standard SWD. We also note that increasing the number of bins (making the quantization finer) does not improve the accuracy of our full approach.
@@
Effect of number of pyramid levels and window size. We ablate the two hyperparameters of the time-series feature extraction: the number of pyramid windows used $L$, and the number of samples per window $\tau$ (see Sec.4.2). We find that in both cases the results are not sensitive to the chosen parameters (Tab.14,15).
@@
Appendix G Using the Central Limit Theorem for Set Anomaly Detection
We model the features of each window $f$ as a normal set as IID observations coming from a probability distribution function $p(f)$. The distribution function is not assumed to be Gaussian. Using a Gaussian density estimator trained on the features of elements observed during training is unlikely to be effective for element-level anomaly detection (due to the non-Gaussian $p(f)$).
An alternative formulation to the one presented in section 3, is that each feature $f$ is multiplied by projection matrix $P$, and then each dimension is discretized and mapped to a one-hot vector according to its relevant histogram bin. This formulation therefore maps the representation of each element to a sparse binary vector. The mean of this one-hot vector representation of elements in the set recovers the normalized histogram descriptor precisely (therefore this formulation is equivalent to the one in section 3). As the histogram is a mean of the one-hot representations of elements, it has superior statistical properties. In particular, the Central Limit Theorem states that under common conditions the sample mean follows the Gaussian distribution. While typically in anomaly detection only a single sample is presented at a time, the situation is different when treating samples as sets. Although the elements are often not IID, given a multitude of elements, an IID approximation may still be applicable. This explains the high effectiveness of Gaussian density estimation in our formulation.
Appendix H Further Discussion
Incorporating deep features for time series data. Our method outperforms the state-of-the-art in time series anomaly detection without using deep neural networks. While this is an interesting and surprising result, we believe that deep features will be incorporated into similar approaches in the future. One direction for doing this is replacing the window projection features with a suitable deep representation, while keeping the set descriptors and Gaussian modeling steps unchanged.
Fine-tuning deep features for anomaly detection. Following recent works in anomaly detection and anomaly segmentation, we used fixed pre-trained features as the backbone of our method. Although some methods fine-tune deep features for anomaly detection based on the normal-only training set, we keep them constant. Doing so allows an interpretable examination of the relative strength of our novel scoring function with respect to prior works that use fixed features. Yet, we expect that fine-tuning such features could lead to further gains in the future.